OpenStudy (anonymous):

classify the discontinuities (if any) for the given function: f(x) = {2, x

OpenStudy (dumbcow):

discontinuous at x=0,1,3

OpenStudy (anonymous):

the answer choices are (a) The function has a removable discontinuity at x = 1 b) The function has a jump discontinuity at x = 0 and 3 c) The function has an infinite discontinuity at x = 3 d) The function is continuous for all x e) The function has a removable discontinuity at x = 3

OpenStudy (anonymous):

would it be answer choice a?

OpenStudy (dumbcow):

no it would def be B) at x=0,1,3 the y value jumps up or down

OpenStudy (anonymous):

thanks, could you also explain how you got that answer?

OpenStudy (dumbcow):

it helps if you try to draw the graph for all neg values y=2 for x between 0 and 1, y=x^2 for x between 1 and 3, y=1 for x>3, y=x At x=0 you have a jump from 2 to 0 because y=x^2=0^2=0 when you have a break in the graph like this it is a jump discontinuity here is site on discontinuities http://en.wikipedia.org/wiki/Jump_discontinuity#Examples

OpenStudy (anonymous):

thanks!